You have 40 cars. There 9 times as many white cars as there are yellow cars. How many white and yellow cars are there
Answer by ReadingBoosters(3246)

x = white cars
y = yellow cars
x+y = 40
9y = x
Substitution
9y + y = 40
10y = 40
10y/10 = 40/10
y = 4 yellow cars
x = 9(4) = 36 white cars